### v1

# Rate, Explain and Cite (REC): Enhanced Explanation and Attribution in Automatic Evaluation by Large Language Models

Abstract
LLMs have demonstrated impressive proficiency in generating coherent and high-quality text, making them valuable across a range of text-generation tasks. However, rigorous evaluation of this generated content is crucial, as ensuring its quality remains a significant challenge due to persistent issues such as factual inaccuracies and hallucinations. This paper introduces two fine-tuned general-purpose LLM autoevaluators, REC-12B and REC-70B—specifically designed to evaluate generated text across several dimensions: faithfulness, instruction following, coherence, and completeness. These models not only provide ratings for these metrics but also offer detailed explanations and verifiable citations, thereby enhancing trust in the content. Moreover, the models support various citation modes, accommodating different requirements for latency and granularity. Extensive evaluations on diverse benchmarks demonstrate that our general-purpose LLM auto-evaluator, REC-70B, outperforms state-of-the-art LLMs, excelling in content evaluation by delivering better quality explanations and citations with minimal bias. It achieves Rank #1 as a generative model on the RewardBench leaderboard [^1] https://huggingface.co/spaces/allenai/reward-bench under the model name TextEval-Llama3.1-70B. Our REC dataset and models are released at https://github.com/adelaidehsu/REC.
Rate, Explain and Cite (REC): Enhanced Explanation and Attribution in Automatic Evaluation by Large Language Models
Aliyah R. Hsu1,2thanks: Work done during an internship at Salesforce., James Zhu1, Zhichao Wang1, Bin Bi1, Shubham Mehrotra1, Shiva K. Pentyala1, Katherine Tan1, Xiang-Bo Mao1, Roshanak Omrani1, Sougata Chaudhuri1, Regunathan Radhakrishnan1, Sitaram Asur1, Claire Na Cheng1, Bin Yu2 1Salesforce AI Platform, 2UC Berkeley Correspondence: aliyahhsu@berkeley.edu, james.zhu@salesforce.com
1 Introduction
Figure 1: Illustration of task prompt, context, response, and model outputs of Rating, Explanation, and Citations (REC). Here we use summarization evaluation as an example.
Large Language Models (LLMs) have demonstrated impressive capabilities in generating high quality coherent text and are deployed in applications for various text-generation tasks (Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023; Touvron et al., 2023). In order for LLMs to provide up-to-date information or to perform knowledge-intensive tasks (Lewis et al., 2020), Retrieval Augmented Generation (RAG) system (Chen et al., 2017; Borgeaud et al., 2021; Izacard et al., 2022; Guu et al., 2020) has been used widely. RAG involves first retrieving document chunks that are relevant to the generation task from an external knowledge source, and performing the generation using an knowledge-augmented prompt. However, it’s of paramount importance to ensure that the generated text is reliable and can be trusted by a human as LLMs often suffer from factual incorrectness and hallucination of knowledge (Ji et al., 2022; Zhang et al., 2023; Shuster et al., 2021). Towards this goal, we propose the generated text needs to be evaluated along various dimensions shown below:
  • Faithfulness: Did the LLM generate factually correct response given the context? (Rating + Explanation task)
  • Instruction Following: Did the LLM generate response follow the instructions provided in the prompt? (Rating + Explanation task)
  • Coherence: Did the LLM generate a coherent response? (Rating + Explanation task)
  • Completeness: Did the LLM generate a complete response including all details? (Rating + Explanation task)
  • Citation: If the LLM generated factually correct response, can we provide evidence for where that response came from? (Citation task)
In this paper, we introduce fine-tuned models for the evaluation tasks listed above that can form the basis for all trust-related evaluations for generative AI applications. The model was fine-tuned to not only provide ratings for the metrics but also explanations + citations for why it rated the generation so. For instance, if the faithfulness metric is rated as “factually correct", the model will output explanation with verifiable citations as to why it rated the generation as factually correct. An example of the model output is shown in Figure 1.
The listed metrics are essentially Rating and Explanation tasks for our fine-tuned models which takes as input: (1) another task LLM’s hydrated prompt (2) a generated response by the task LLM. The output would be a rating and an associated explanation for that rating as shown in Figure 2.
Figure 2: Model Inputs and Outputs when providing Rating with Explanations and citations.
The Citation task takes as input (1) a provided context (2) a generated response. The output would be a citation tagged response that provides evidence for parts of generated text as shown in Figure 2.
To differentiate from the content quality citations  [^2] For content quality citation task, we cite from another task LLM’s hydrated prompt (context) and include the citations with the generated explanation (response) by the LLM autoevaluator. we obtain when using our model as an autoevaluator for verifiable evaluation by performing Rating, Explanation and Citation, we emphasize when performing the Citation task alone, our model is able to conduct general RAG citations, where we cite from the retrieved chunks (context) and tag the citations with the generated answer (response). For both kinds, we support various citation modes as listed below (See Appendix A for examples on RAG citation modes):
  • Post-fix citation: If the generation is few sentences long, we place the citation at the end of the response and refer to the context id.
  • Post-fix citation mode with snippet: If the generation is few sentences long, we place the citation at the end of the response and refer to the exact sentence within that context that supports the generation.
  • Inline citation: If the generation is long, we place the citation at appropriate locations within the generation and refer to the context id.
  • Inline citation mode with snippet: If the generation is long, we place the citation at appropriate locations within the generation and refer to the exact sentence within that context that supports the generation.
These different modes of citations were designed to cater to different trade-offs between latency and granularity of citation. For instance, the Post-fix citation mode would be the fastest as there is no need for our fine-tuned model to generate claims or snippet. It simply has to generate the reference to the cited context. On the other hand, the inline citation mode with snippet has to place the citation inline with the generated response and also point to a snippet within the corresponding cited context. This mode is the most granular but can also increase latency as it needs to generate more output tokens.
Our model is the first to enable citation in both automatic evaluation and general task output with scalability and efficiency, while prior work often solely focus on providing rating and explanation in automatic evaluation (Zhu et al., 2023; Li et al., 2023a; Wang et al., 2023b; Vu et al., 2024a), or requires iterative prompting (Sun et al., 2023; Cao and Wang, 2024; Ye et al., 2023) or human annotation (Malaviya et al., 2024) for general citations generation. Our contributions include the following:
  • A novel general-purpose LLM autoevaluator that comes in two sizes: 12B and 70B, that can generate better quality Rating, Explanation and Citations (REC), with little to no trade-off in general instruction task performance evaluated on various public benchmark datasets and our own dataset described below.
  • A curated dataset for citations and explanations fine-tuning to facilitate future research, which is the first public dataset containing both content quality citations and RAG citations.
  • A single model that can perform different modes of citation to cover the trade-off between latency and granularity of citation. Such model with generalized capability largely simplify the deployment complexity in production.
We provide extensive evaluations of our REC models on diverse benchmarks evaluating LLMs’ (1) RAG citation capability: ALCE (Gao et al., 2023), and ExpertQA (Malaviya et al., 2024); (2) content quality citation capability: human evaluation on ABCD summarization (Chen et al., 2021); (3) general capabilities: RewardBench (Lambert et al., 2024), and LLM-AggreFact (Tang et al., 2024a); (4) cognitive bias: CoBBLEr (Koo et al., 2024), against 7 state-of-the-art (SOTA) LLMs, such as GPTs and Claude. We show our REC models outperforms the state-of-the-art baselines on these benchmarks, and that the baselines often struggle to provide accurate citations, undermining their reliability. Specifically, our REC-70B model achieves Rank #1 as a generative model on the RewardBench leaderboard under the model name TextEval-Llama3.1-70B. The model outperforms the second ranked model in three dimensions: Chat Hard, Safety, Reasoning, which are key capabilities required to generate accurate explanation and citation outputs.
2 Related Work
Verifiable Text Generation
Providing citations to attributable information source (Liu et al., 2023a; Gao et al., 2023) has been proposed as one of the approaches to mitigate hallucination and improve the factuality of LLM generation. Such work falls into the general RAG citation setting defined in our work, where the aim is to generate verifiable content through citing from the provided document chunks. Initial efforts fine-tune LLMs using either human-written examples (Nakano et al., 2021) or machine-generated examples verified by human (Menick et al., 2022), but their privately maintained training data limits further research. With the advent of more capable LLMs (OpenAI, 2023; Jiang et al., 2023), most existing work relies on zero-shot prompting or few-shot prompting to cite document chunks during generation (Kamalloo et al., 2023; Gao et al., 2023; Liu et al., 2023a), although the quality of their generated citations leaves large room for improvement (Malaviya et al., 2024). Other work utilizes an additional natural language inference (NLI) model to add citations (Gao et al., 2022; Chen et al., 2023). To improve upon prior work, we propose a fine-tuned LLM that can generate better-grounded responses supported with citations of various modes, with the fine-tuning dataset released to the public to facilitate future research and no additional NLI models needed.
General-purpose Autoevaluators
Collecting human annotations to evaluate LLM is not only costly, time-consuming, but hard to replicate (Ouyang et al., 2022; Zheng et al., 2023; Chiang and Lee, 2023), as a result LLMs become a natural automated proxy to evaluate LLM capabilities on various benchmarks (Bai et al., 2024; Bubeck et al., 2023; Chiang et al., 2023; Fu et al., 2023; Liu et al., 2023b; Wang et al., 2023a). Existing LLM autoevaluators often judge LLM outputs by expressing “preference" over outputs from a reference model (Dubois et al., 2024; Li et al., 2023b; Yuchen Lin et al., 2024), or by providing rating and explanation according to a user-defined metric as a direct assessment (Li et al., 2023a; Wang et al., 2023b, 2024b). Closer to our work, Jiang et al. (2024) and Xu et al. (2023) fine-tunes LLMs to generate rating, explanation, and a detailed analysis to pinpoint errors in the response evaluated. Unlike prior work, our fine-tuned model is a more generalizable autoevaluator since it provides content quality citations for both good and bad evaluated responses, in addition to the rating and explanation. This allows users to not only be able to diagnose where errors are from but provide supporting evidence when a generation is good, which are both important to establish trust in the model.
3 Data Collection
To fine-tune our general-purpose LLM autoevaluator, we meticulously collect a multitask mixture of data encompassing a broad spectrum of LLM capabilities (Section 3.1), and we denote our collected dataset as REC-Data. Due to a lack of publicly available content quality citation datasets, we leverage synthetic data generation using LLM to curate such data (Section 3.2) to facilitate the research community. We perform rule-based automatic quality check followed by a unified task formatting to post-process our data for instruction fine-tuning (Section 3.3).
3.1 Task Types
To enhance the explainability in the automatic evaluation of our LLM autoevaluator, while maintaining its general instruction-following capability, we gather datasets from a diverse range of task types (See detailed REC-Data distribution in Appendix B). These include:
  • Pairwise Evaluation: Compare two responses at the same time and express a preference according to evaluation criteria.
  • Pointwise Evaluation: Evaluate specific aspects of a response independently according to evaluation criteria and provide a rating.
  • Open-ended Evaluation: Evaluate a response independently and provide a free-form explanation, usually to support either pairwise or pointwise evaluation.
  • Citation: Evaluate a response alongside the context, and provide citations for verifiable answer attribution.
  • General Instruction: Generate a response as instructed (no evaluation tasks in this type), such as summarization and QA.
3.2 Synthetic data generation
We leverage synthetic data generation for some of the tasks, including pointwise evaluation and citation. For synthetic pointwise evaluation data, we leverage Mistral-7B to rate a response according to one of the 4 metrics listed in Section 1. For synthetic citation data generation, due to the challenging nature of citation tasks, we leverage a more powerful instruction-tuned LLM based on Llama-3.1-70B for both RAG and content quality citations. Example prompts used to generate the synthetic data is provide in Appendix C.
3.3 Post-processing and Unified Task format
Knowing that LLM-generated outputs is susceptible to hallucination, we perform automatic rule-based content quality check to ensure the quality of the synthetic data. During post-processing, we filter invalid synthetic LLM outputs that have either invalid output formats, or non-factual content cited by checking exact-match of the citation with the context. Our final REC-Data has in total around 140k datapoints, each containing a task prompt and a completion.
ASQA QAMPARI ELI5 AVG
Fluency Correct Citation F1 Correct Citation F1 Fluency Correct Citation F1
Mistral-7B 86.55 19.83 65.31 15.65 16.24 42.43 9.89 47.28 37.90
Mistral-Nemo 80.42 21.49 67.49 17.11 20.85 42.29 15.49 35.61 37.59
Llama-3.1-70B 77.95 25.51 67.07 17.72 16.90 66.07 14.06 36.56 40.23
Claude-3-Opus 89.22 24.08 69.29 0.54 1.92 45.84 11.72 34.69 34.66
GPT-3.5 86.06 23.63 81.02 21.20 19.79 16.35 12.23 51.80 39.01
GPT-4 68.99 25.59 71.73 20.48 25.63 46.98 14.19 43.92 39.69
GPT-4o 79.87 28.73 67.29 20.97 27.11 32.71 17.70 27.11 37.69
REC-12B 81.85 24.36 70.59 18.18 22.71 48.02 16.27 37.60 39.95
REC-70B 79.35 26.06 68.93 18.59 17.48 66.80 16.92 38.79 41.62
Table 1: Fluency, citation quality and correctness on ASQA, QAMPARI, and ELI5 datasets of ALCE
AutoAIS FActscore
Mistral-7B 26.39 84.03
Mistral-Nemo 41.39 85.78
Llama-3.1-70B 52.11 86.42
Claude-3-Opus 59.84 87.16
GPT-3.5 57.65 86.27
GPT-4 60.18 87.31
GPT-4o 60.86 87.03
REC-12B 62.90 88.22
REC-70B 66.52 89.05
Table 2: General RAG citation quality and answer correctness results on ExpertQA.
4 Model Training
The general-purpose LLM autoevaluator that we trained is available in two sizes, and they are denoted as REC-12B and REC-70B separately. REC-12B is instruction fine-tuned from Mistral-Nemo [^3] https://mistral.ai/news/mistral-nemo, and REC-70B is instruction fine-tuned from Llama-3.1-70B Team (2024). We adopted supervised fine-tuning (SFT) to optimize the models. To accommodate GPU memory constraints, we made the following design choices: we filtered out examples where the combined length of the prompt and the response exceeded 6,144 tokens. We used Low-Rank Adaptation (LoRA) during fine-tuning Hu et al. (2021), with a rank of $r=256$ for REC-12B and $r=64$ for REC-70B. Additionally, We applied 4-bit quantization for REC-70B during initialization. Both models were trained using a learning rate of $1\times 10^{-4}$, with data shuffled and for a single epoch. REC-12B was trained with a batch size of 2 and a gradient accumulation factor of 8, whereas for REC-70B we used a batch size of 1 with the same gradient accumulation factor. Both models were trained on eight NVIDIA H100 GPUs, each with 80GB of memory. The total training time was approximately 4.5 hours for REC-12B and 5.5 hours for REC-70B.
5 Evaluation
In this section, we evaluate our REC models on diverse benchmarks assessing LLMs’ (1) RAG citation capability: ALCE (Gao et al., 2023), and ExpertQA (Malaviya et al., 2024); (2) content quality citation capability: human evaluation on ABCD summarization (Chen et al., 2021); (3) general capabilities: RewardBench (Lambert et al., 2024), and LLM-AggreFact (Tang et al., 2024a); (4) cognitive bias: CoBBLEr (Koo et al., 2024), each of which is an important measure to a general-purpose LLM autoevaluator. We compare our results against 7 SOTA LLMs, including Misral-7B (Mistral-7B-Instruct-v0.2), Mistral-Nemo (Mistral-Nemo-Instruct-2407), Llama-3.1-70B (Llama-3.1-70B), Claude-3-Opus (claude-3-opus-20240229), GPT-3.5 (gpt-3.5-turbo), GPT-4 (gpt-4-turbo), GPT-4o (gpt-4o).
5.1 RAG citation & Correctness
ALCE
The ALCE (Automatic LLMs’ Citation Evaluation) benchmark is designed to assess the ability of LLMs to generate text with accurate and relevant citations. ALCE addresses this by providing a framework that evaluates the quality of citations in text generated by LLMs, focusing on three key dimensions: fluency, correctness, and citation quality. The ALCE benchmark is built on three datasets: ASQA (a short-answer question dataset), QAMPARI (which provides lists of correct answers), and ELI5 (a long-form question-answer dataset). For each dataset, ALCE evaluates how well the model generates text that is not only fluent and accurate but also properly supported by citations. The evaluation includes metrics like citation precision (ensuring all cited sources are relevant) and citation recall (ensuring all necessary sources are cited). ALCE does not provide training data but instead measures citation performance through retrieval-based systems that simulate real-world information-seeking tasks, where the model retrieves relevant passages and uses them to support its answers.
As shown in Table 1, our REC models rank highly in terms of overall average performance, with REC-70B achieving an average score of 41.62, the highest among compared models, and REC-12B reaching 39.95. This highlights the robustness and balance of these models across fluency, correctness, and citation quality. REC offers a well-rounded solution across different types of questions, making it highly reliable for real-world applications where both citation quality and answer correctness are essential. Specifically, while GPT-4 excels in citation quality, the much smaller REC models outperform all the other competitors.
Instruction-following Completeness Faithfulness
Rating accuracy Explanation accuracy Citation (F1) Rating accuracy Explanation accuracy Citation (F1) Rating accuracy Explanation accuracy Citation (F1)
Mistral-7B 1.00 1.00 0.05 0.92 0.97 0.04 0.97 0.96 0.08
Mistral-Nemo 0.96 0.96 0.51 0.93 0.95 0.11 0.95 0.95 0.29
Llama-3.1-70B 0.98 0.98 0.73 0.96 0.97 0.54 0.97 0.96 0.37
Claude3-Opus 1.00 1.00 0.67 0.98 0.99 0.46 0.98 0.97 0.59
GPT-3.5 0.88 0.99 0.48 0.88 0.58 0.21 0.78 0.99 0.29
GPT-4 0.97 0.97 0.38 0.96 0.96 0.38 0.96 0.96 0.28
GPT-4o 0.95 0.95 0.35 0.93 0.93 0.56 0.96 0.71 0.59
REC-12B 1.00 1.00 0.50 0.94 0.95 0.62 0.94 0.98 0.60
REC-70B 1.00 1.00 0.50 0.96 0.96 0.67 0.97 0.98 0.63
Table 3: Human evaluation results on ABCD summarization content quality evaluation task. For the correctness of rating and explanation, we report the average results of the two labelers. For citations, we evaluate the intersection of the results from both labelers against the machine-generated citations to compute F1.
ExpertQA
ExpertQA Malaviya et al. (2024) is designed to evaluate LLMs’ ability to generate accurate and well-attributed responses in technical and high-stakes fields, such as medicine and law. It was developed by involving experts from 32 different fields, who contributed 2,177 domain-specific questions based on their knowledge. In total, 484 participants helped curate these questions. LLMs were then used to generate responses to the questions, followed by human experts evaluating the quality of these responses on several criteria, including factual correctness, completeness of attribution, source reliability, and informativeness. We adopt the metrics used in its paper: AutoAIS Gao et al. (2022), which is similar to citation recall, and FActscore Min et al. (2023) which measures the percentage of generated claims that are factual. We generate claims from gold long-form answers with every compared model, and use an NLI model Honovich et al. (2022) to determine the attribution of claim-evidence pairs in the ExpertQA dataset.
As shown by the zero-shot results in Table 2, our REC-12B and REC-70B achieve the highest AutoAIS and FActscore scores, which not only indicates their best citation quality but also the generalizability of their attribution capability across domains.
5.2 Rating & Explanation & Citation
ABCD
To evaluate our LLM autoevaluator’s citation capability more comprehensively, besides examining its general RAG citation capability on QA tasks as reported in Section 5.1, we evaluate the content quality citation capability on a summarization task, the ABCD dataset Chen et al. (2021), in this section.
The ABCD dataset Chen et al. (2021) contains customer support conversation transcripts. We generate summaries by prompting Mistral-7B (Jiang et al., 2023) with: “Summarize knowledge from transcripts after they’ve ended, including the customer issue and resolution." to form (dialogue, summary) pairs for evaluation. LLM autoevaluators then take as input the summarization task prompt hydrated with the dialogue, and the summary, to generate rating, explanation, and citations according to metrics defined in Section 1. Note that we use a held-out set of 50 examples in this experiment to make sure there exists no data overlap with the ABCD subsets used during training. However, unlike all the other evaluations included in this paper, this evaluation is in-distribution.
Different from the general RAG citation benchmarks evaluated in Section 5.1, which contain human annotated citation groundtruth, there exists no well-established benchmarks to evaluate content quality citations. Hence, we seek human annotations to serve as a reference standard in this task. The final citation outputs were labeled by 12 machine learning experts where each has a graduate degree in computer science or related fields. They were asked to complete three tasks: rating correctness evaluation, explanation correctness evaluation, and to provide manually-written citations for the claims in explanation (See Appendix D for human labeling guideline).
We assign two labelers for each model and find the average inter-rater agreement is around 95%. For the correctness of rating and explanation, we report average results of the two labels, while for citations, we evaluate the intersection of the results of the two labelers against the machine-generated citations to compute F1 score as shown in Table 3. Due to the high quality summary generated by the summarizer LLM, most LLM autoevaluators obtain high rating and explanation accuracy. As for the citation task, both REC-12B and REC-70B models consistently outperform other models in F1 score.
Chat Chat Hard Safety Reasoning AVERAGE
Mistral-7B 76.10 54.30 81.50 53.90 66.50
Mistral-Nemo 86.60 59.60 74.80 70.30 72.80
Llama-3.1-70B 97.60 58.90 73.00 78.50 77.00
Claude-3-Opus 94.70 60.30 86.60 78.70 80.10
GPT-3.5 92.20 44.50 65.50 59.10 65.30
GPT-4 95.30 74.30 87.60 86.90 86.00
GPT-4o 96.10 76.10 88.10 86.60 86.70
REC-12B 91.90 86.60 92.00 93.60 91.00
REC-70B 94.10 90.10 93.20 96.40 93.50
Table 4: General LLM capabilities evaluation on RewardBench. Scores for all the compared LLMs are obtained from the RewardBench Leaderboard Lambert et al. (2024).
LLM-FactVerify Wiki-FactVerify Summarization Long-form QA AVERAGE
Mistral-7B 68.04 56.82 55.80 62.80 60.86
Mistral-Nemo 70.18 65.27 62.14 60.88 64.62
Llama-3.1-70B 75.25 68.84 69.56 76.98 72.66
Claude-3-Opus 76.82 76.52 70.65 76.82 75.20
GPT-3.5 72.81 69.53 68.60 70.40 70.33
GPT-4 77.44 77.22 72.45 76.92 76.01
GPT-4o 76.31 79.48 72.18 75.22 75.80
REC-12B 78.78 85.34 72.51 77.23 78.46
REC-70B 79.67 87.48 74.04 79.12 80.07
Table 5: LLM-AggreFact performance across four common use-cases: LLM-FactVerify (ClaimVerify + FactCheck + Reveal), Wiki-FactVerify (WiCE), Summarization (AggreFact + TofuEval), and Long-form QA (ExpertQA + LFQA). REC-70B outperforms all tested LLM-as-a-Judge models in all four categories.
Order Bandwagon Compassion Selective Salience Distraction Frequency AVERAGE
Mistral-7B 0.163 0.518 0.346 0.671 0.603 0.0143 0.1406 0.3508
Mistral-Nemo 0.126 0.543 0.238 0.725 0.595 0.0196 0.0059 0.3218
Llama-3.1-70B 0.115 0.251 0.318 0.545 0.599 0.0116 0.0705 0.2729
Claude-3-Opus 0.078 0.081 0.380 0.483 0.589 0.0093 0.0087 0.2327
GPT-3.5 0.164 0.877 0.371 0.851 0.644 0.0029 0.0183 0.4183
GPT-4 0.092 0.098 0.399 0.431 0.569 0.0015 0.0046 0.2279
GPT-4o 0.075 0.086 0.417 0.468 0.578 0.0089 0.0112 0.2349
REC-12B 0.089 0.225 0.291 0.526 0.583 0.0108 0.0048 0.2471
REC-70B 0.073 0.095 0.258 0.487 0.572 0.0091 0.0043 0.2141
Table 6: Bias evaluation on CoBBLEr (lower is better).
5.3 General Capabilities
RewardBench
The RewardBench dataset assesses LLMs’ performance across several critical abilities using curated chosen-rejected response pairs Lambert et al. (2024). The evaluation process involves determining the model’s win percentage based on its ability to score the “chosen" response higher than the “rejected" one. Specifically, it’s focused on testing chat performance, safety measures, and reasoning (both in terms of code and math skills), while controlling for potential biases such as overfitting to prior datasets.
First we create prompts following the RewardBench evaluation format, where each prompt consisting of a query and two potential responses. The task for LLM is to evaluate these responses and determine which one is better. The performance of the reward models is quantified by calculating the win percentage for chosen-rejected pairs associated with each prompt. A “win" is defined by a scenario where the LLM presents a preference for the selected answer. This quantitative measure provides a clear benchmark for evaluating model performance across tasks. By averaging the results across the 4 core sections: Chat, Chat Hard, Safety, and Reasoning, a comprehensive measure of each model’s win percentage is derived. Our REC-12B achieves SOTA performance among all generative LLM with fewer than 20 billion parameters on the RewardBench leaderboard by October 15th, 2024, while our REC-70B achieves SOTA among all generative LLMs within the RewardBench Leaderboard by October 15th, 2024 as shown in Table 4.
LLM-AggreFact
This is a benchmark for measuring the grounding capabilities of autoevaluators. Given a reference document and a claim, the autoevaluator determines if the claim is fully supported by the document. This holistic benchmark combines 10 attribution datasets used in recent studies on LLM factuality. Table 5 presents our attribution results on LLM-AggreFact Tang et al. (2024b), categorized into four common use-cases: (1) LLM-FactVerify: fact verification of LLM-generated responses, (2) Wiki-FactVerify: evaluating correctness of Wikipedia claims, (3) Summarization: assessing faithfulness of summaries, and (4) Long-form QA: evaluating long-form answers to questions. Our proposed model REC-70B outperforms all other models in all four categories. REC-70B achieves the highest overall average performance of 80.07, while the next-best baseline model REC-12B obtains a score of 78.46. In long-form QA attribution evaluation, our REC-70B model outperforms GPT-4 (80.07 vs 76.01), demonstrating its strong performance across all categories.
Given some of the models like GPT-3.5 have lower context windows, we removed examples longer than 16K tiktoken tokens from our evaluation. we use the same prompt instructions across categories, as shown in Appendix E. To reduce model API costs, we randomly sample 256 examples per evaluation task similar to the approach used in the FLAMe Vu et al. (2024b).
5.4 Bias Testing
CoBBLEr
Recent studies have found that LLM-as-a-Judge often exhibits cognitive biases, such as preferences for verbosity, egocentrism, bandwagon, and an overly authoritative tone Wang et al. (2024a); Koo et al. (2024); Zheng et al. (2024); Chen et al. (2024). To investigate the biases of the compared models, we evaluate them on the CoBBLEr benchmark (Cognitive Bias Benchmark for LLMs as EvaluatoRs) Koo et al. (2024). This dataset is designed to evaluate the quality and reliability of LLMs when used as automated evaluators in a question-answering (QA) setting. It assesses the presence of six cognitive biases, both implicit and induced, when LLMs are tasked with ranking responses generated by various other models. CoBBLEr’s core objective is to identify the extent of bias in LLM evaluation outputs.
The CoBBLEr dataset includes 50 QA instructions, randomly selected from two well-established benchmarks: BIG-bench bench authors (2023) and ELI5 Fan et al. (2019). 16 LLMs, both open and closed-source models, generate responses to these instructions. The evaluations involve pairwise comparisons between the responses of two models, wherein each model also acts as an evaluator to rank its own and others’ outputs. The biases tested are categorized into two groups: (1) Implicit biases, such as egocentric bias (where a model tends to prefer its own outputs), and (2) Induced biases, such as order bias, where the ranking of responses is influenced by their order in the evaluation.
Table 6 presents the performance of the compared models on the CoBBLEr benchmark with metrics of order bias, bandwagon effect, compassion, selective bias, salience, distraction, and frequency. Note that lower scores indicates better performance  (i.e., fewer biases) on CoBBLEr. Overall, our REC-70B model has the best average performance with a score of 0.2141, followed closely by GPT-4 (0.2279) and GPT-4o (0.2349). REC-12B, on the other hand, outperforms all the other same-sized and smaller models (Mistral-Nemo and Mistral-7B), and even some larger models (Llama-3-70B and GPT-3.5). The results suggests that training with REC leads to fewer cognitive biases and more consistent evaluation capabilities. In contrast, off-the-shelf LLMs can be more influenced by factors such as the order of responses or the length of the text.
6 Conclusions
LLMs excel in generating coherent and high-quality text for various tasks, but evaluating content quality is challenging due to issues like factual inaccuracies and hallucinations. This paper introduces two fine-tuned models—REC-12B and REC-70B—designed to evaluate generated text across several dimensions: faithfulness, instruction following, coherence, completeness. These models not only provide ratings for these metrics but also offer detailed explanations and verifiable citations, i.e., Ratings, Explanations, and Citations (REC), thereby enhancing trust in the content. To achieve this goal, a curated dataset for explanations and citations fine-tuning is proposed to facilitate future research, which is the first public dataset containing both content quality citations and RAG citations. The models support various citation modes, including Post-fix citation, Post-fix citation mode with snippet, Inline citation, and Inline citation mode with snippet to balance the trade-off between latency and granularity of citation. Extensive evaluations on diverse benchmarks—ALCE, ExpertQA, ABCD summarization, RewardBench, LLM-AggreFact, and CoBBLEr—demonstrate that the fine-tuned LLM autoevaluator, REC-70B, outperforms state-of-the-art models. It excels in content evaluation, delivering clear explanations and ensuring citation accuracy.
Limitations
Our work introduces the innovative Rate, Explain, and Cite (REC) model for trusted automated AI evaluations, offering significant benefits but with a few limitations that future research will address. First, our work is limited to the baselines we compare, as non-LLM citation methods, such as semantic similarity, were not included. Including such techniques could serve as valuable baselines to further validate our model’s generalization and effectiveness. Second, the need for post-generation explanations and citations may introduce latency. Future work will compare latencies across models and explore optimizations to improve speed. We are also considering fine-tuning the snippet output using starting and ending character numbers to optimize for latency. Finally, even though the 12B base model is capable of generating multi-lingual output, our current evaluations are limited to English datasets. Expanding to multilingual evaluations is a key goal for future research.
Ethical Considerations and Risks
Our contributions present a novel general-purpose autoevaluator that supports various modes of citations, as well as a dataset with citations and explanations to support further research. There are a few ethical considerations for users of such technology. As with the use of fine-tuning and RAG, it is important to be aware of potential propagating biases and feedback loops. Biased data used for training and/or retrieval could lead to model outputs that reinforces such biases, impacting model quality and fairness. Although we extensively evaluated our LLM autoevaluator and found it to be safe, unbiased, and factual, users should consider whether such considerations hold true for the input source itself. Of note, even established sources such as knowledge banks may contain inaccurate or inconsistent information. Finally, though we aimed to be transparent and consistent for our data labeling tasks, using clear instructions and validations, note that data labeling tasks are inherently subject to the bias of labeler background and experiences.
Acknowledgments
We appreciate Yilun Zhou, Shafiq Rayhan Joty, Peifeng Wang, and Austin Xu from Salesforce AI research for insightful discussions and help to benchmark model performance. We gratefully acknowledge partial support from NSF grant DMS-2413265, NSF grant 2023505 on Collaborative Research: Foundations of Data Science Institute (FODSI), the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and 814639, NSF grant MC2378 to the Institute for Artificial CyberThreat Intelligence and OperatioN (ACTION), and a Berkeley Deep Drive (BDD) grant from BAIR and a Dean’s fund from CoE, both at UC Berkeley.
Appendix A Various RAG Citation Modes that the REC Models Support
Here are example output templates of the different citation modes that our model supports and has been evaluated on. Users can easily process information from the JSON objects into the corresponding natural language format they desire. For all of these output styles, here’s the example response and context that we’ve used to feed into the model.
  • Retrieved Context:
    ID 1233
    Photosynthesis is the process by which green plants and some other organisms utilize sunlight to synthesize their food. This remarkable process involves the conversion of carbon dioxide and water into glucose and oxygen, facilitated by chlorophyll. It serves as the foundation for energy production in these organisms and plays a crucial role in maintaining the balance of oxygen in our atmosphere.
    ID 1422
    Cellular respiration refers to a series of metabolic reactions that occur within the cells of organisms. This process transforms biochemical energy from nutrients into adenosine triphosphate (ATP), which is essential for cellular functions. During cellular respiration, waste products are released, and oxygen is a vital element needed for the process to function efficiently. This energy conversion is critical for sustaining life.
    ID 4431
    DNA replication is the mechanism by which a cell duplicates its DNA in preparation for cell division. This process ensures that genetic information is accurately copied and passed on to daughter cells. The enzyme DNA polymerase is instrumental in DNA replication, playing a crucial role in maintaining genetic fidelity and continuity across generations.
  • LLM Generated Response to Cite:
    Photosynthesis is a process that converts carbon dioxide and water into glucose and oxygen using sunlight and chlorophyll. This process occurs in green plants and certain other organisms.
A.1 Post-fix citation
{  "citations": [    {      "context_id": "1233"    }  ]}
A.2 Post-fix citation with snippet extracted verbatim from retrieved context
{  "citations": [    {      "context_id": "1233",         "snippet": "Photosynthesis is the         process by which green plants and         some other organisms utilize         sunlight to synthesize         their food. This remarkable         process involves the         conversion of carbon         dioxide and water into         glucose and oxygen,         facilitated by chlorophyll"    }  ]}
A.3 Inline citation with claims extracted verbatim from LLM generated response
{  "citations": [    {      "context_id": "1233",         "claim": "Photosynthesis is a process         that converts carbon dioxide         and water into glucose and         oxygen using sunlight and         chlorophyll."    }  ]}
A.4 Inline citation with snippet
{  "citations": [    {      "context_id": "1233",       "claim": "Photosynthesis is a process         that converts carbon dioxide         and water into glucose and         oxygen using sunlight and         chlorophyll.",       "snippet": "Photosynthesis is the         process by which green         plants and some other         organisms utilize sunlight         to synthesize their food. This         remarkable process involves         the conversion of carbon         dioxide and water into         glucose and oxygen,         facilitated by chlorophyll"    }  ]}
A.5 Prompt for Inline Citation
Below is the prompt that we used to generate inline citations. This can be modified to generate responses in each of the different RAG citation modes shown in the sections above.
Your task is to providecitations for a generated responsefrom an LLM for a RAG(Retrieval Augmented Generation)application. You will be provided twosections below. The RETRIEVED CHUNKSsection is the context fromwhere you will generate citations.The LLM GENERATED ANSWER sectionis for the LLM generated answerfrom where you will generate claims.###Respond using the following JSON format:Desired format:{   "citations":[      {         "context_id":"<context_id>",         "claim": "<claims identified from         LLM GENERATED ANSWER>"      }   ]}###INSTRUCTIONSIn order to generate citations,follow these steps below:1. Review the "LLM GENERATED ANSWER"section below and extract a set of claims,VERBATIM and set the "claim" in the outputjson.2. Your task is to generate  a citationfor each of the claims identifiedin Step 1, by following the steps below:    i. For each claim, identify the    most relevant chunk and the corresponding    "context_id" from the "RETRIEVED CHUNKS"    section that supports the claim.    ii. For the identified chunk,    set the "context_id" in the output    to the corresponding context_id.    Set "context_id" to "None" if none    of the chunks in the "RETRIEVED CHUNKS"   section supports the claim.3. Ensure that you structure yourresponse to adhere to the desiredoutput JSON format.###RETRIEVED CHUNKS:{retrieved_chunks}###LLM GENERATED ANSWER:{answer}###Response (JSON only):
Appendix B REC-Data Details
For general instruction tasks, we include summarization on ABCD (Chen et al., 2021), a dataset containing customer support dialogue transcripts, in five languages. For pointwise evaluation, we leverage HelpSteer2 (Wang et al., 2024c) and perform synthetic data generation (Section 3.2) on the (dialogue, summary) pairs from ABCD, and QA pairs from FloDial. For pairwise and open-ended evaluation tasks, we use Auto-j (Li et al., 2023a), Shepherd (Wang et al., 2023b), Skywork  [^4] https://huggingface.co/datasets/Skywork/Skywork-Reward-Preference-80K-v0.1, HelpSteer2555https://huggingface.co/datasets/nvidia/HelpSteer2, OffsetBias [^6] https://huggingface.co/datasets/NCSOFT/offsetbias, and Code Preference [^7] https://huggingface.co/datasets/Vezora/Code-Preference-Pairs. For citations, we perform synthetic data generation (Section 3.2) on FloDial for both RAG citations and content quality citations, and on ABCD, Auto-j, Shepherd, and Email Thread Summary  [^8] https://www.kaggle.com/datasets/marawanxmamdouh/email-thread-summary-dataset/data for content quality citations.
Note that it’s possible to derive different tasks from the same dataset when different task prompts are assigned to an LLM. We provide a detailed breakdown of REC-Data in Figure 3.
Figure 3: Detailed breakdown of REC-Data distribution.
Appendix C Synthetic Data Generation Prompt Example
C.1 Pointwise Evaluation ("Faithfulness metric")
You will be given a source text given to anotherLLM, and that LLM’s answer.Your task is to rate that LLM’s answeron one quality metric.Please make sure you read and understandthese instructions carefully. Pleasekeep this document open while reviewing,and refer to it as needed.Evaluation criteria:metric_name = ’Faithfulness’metric_scale = ’Yes/No’metric_description = ""The generated answer only contains truthfulcontent, and does not contain inventedor misleading facts that arenot supported by the context.""Evaluation steps:1. Read the source text.2. Read the other LLM’s answer and compareit to the source text. Check if theevaluation criteria is met.3. If the evaluation criteria is met,answer Yes; if the evaluation criteriais not met, answer No.4. Create a JSON response{    "metriclabel": <Yes|No only>,    "justification":    <explanation for the score>}Example:Source text given to another LLM:{query_with_context}The other LLM’s answer:{answer}### Response(JSON Only):
C.2 Rating, Explanation, and Citation Evaluation Prompt
You will be asked to providefeedback on one quality metric for a taskprompt given to another LLM and thatLLM’s output. Given the task prompt as acontext, your job is to give feedback onwhether there are any errors in the output,based on the quality metric. If thereare no errors in the output, pleasealso provide your feedback that there areno errors. Remember your main goal is toprovide feedback so that another LLM canuse your feedback to improve its generation.You need to support statements in yourfeedback that can be linked back to ANOTHERLLM’S TASK PROMPT with citations.Please make sure you read and understandthese instructions carefully. Please keepthis document open while reviewing, andrefer to it as needed.Please only generate your answer followingthe Output JSON format. Do not generateanything else.Evaluation criteria:{metric_name} ({metric_scale})- {metric_description}Respond using the following JSON format basedon the provided ANOTHER LLM’S TASK PROMPT andANOTHER LLM’S OUTPUT:Desired format:{   "answer":<answer>,   "feedback":<generated feedback>,   "statements":[   {       "statement_string":<one statement       extracted from feedback>,       "citations":[       {            "snippet":<snippet>       }]    }]}### INSTRUCTIONS:Here are the instructions to follow togenerate the response.1. Generate a feedback based on ANOTHERLLM’S TASK PROMPT and ANOTHER LLM’SOUTPUT. Explain whether the evaluationcriteria is met.    - Set "answer" to "Yes" if the evaluation    criteria is met, and "No" otherwise.    - Set "feedback" to the    generated feedback.2. For each statement in your feedback,extract relevant sentences that support thestatement from ANOTHER LLM’S TASK PROMPT.3. For each extracted sentence:    - Set "snippet" to be the extracted    sentence.4. Compile relevant snippets for thestatement:    - Set "statement_string" to be    a statement in the generated feedback.    - Set "citations" to be a list of JSON    objects containing all relevant snippets    (from step 3) supporting the    "statement_string".5. Compile the statements into a list ofJSON objects.    - Set "statements" to be a list of    JSON objects containing metadata of    all statements (from step 4) in the    feedback.[BEGIN DATA]### ANOTHER LLM’S TASK PROMPT:Summarize knowledge from transcripts afterthey’ve ended. Summarizing the key points ofthe conversation including customerissue and resolution.### Conversation:{convo}### ANOTHER LLM’S OUTPUT:{answer}[END DATA]### Response(JSON only):
Appendix D Instruction for Claim + Citations Labeling
We’ve asked an LLM to evaluate another LLM’s ’Answer’ for a summarization task given a ’Query’ and ’Context’. The evaluator LLM has generated explanations for its evaluation with supporting citations, as shown in the ’metric_full_explain’, ’Claim’ (parts extracted from the explanation) and ’Citations’ columns in a file separately.
D.1 Instructions
You have two tasks to complete:
  1. 1.
    Evaluate the quality of the ’Claim’ generated by the evaluator LLM based on its corresponding ’Metric’. For each claim, first identify which metric it’s trying to explain against by looking at the ’Metric’ column, and refer to the corresponding label guideline for that metric below. Fill in your answer in ’Claim_correct’. Note, you don’t need to read the ’metric_full_explain’ column to decide if the claim is correct or not.
    • Faithfulness:
      • Choose ’Yes’: A claim for faithfulness is correct if it follows ALL the three criteria: (1) it is a correct judgement, (2) it correctly justifies that the answer either contains only accurate information mentioned in the context or points out the invented facts, and (3) it doesn’t contain irrelevant or extra information in addition to the justification itself.
      • Choose ’No’: otherwise.
    • Instruction-following:
      • Choose ’Yes’: A claim for instruction-following is correct if it follows ALL the three criteria: (1) it is a correct judgement, (2) it correctly justifies why (or why not) the answer is relevant to the query and follows what’s asked in the query, not performing any extra tasks or repeating the same task more than once, and (3) it doesn’t contain irrelevant or extra information in addition to the justification itself.
      • Choose ’No’: otherwise.
    • Coherence:
      • Choose ’Yes’: A claim for coherence is correct if it follows ALL the three criteria: (1) it is a correct judgement, (2) it correctly justifies why (or why not) the answer is easy to understand and free of grammar errors, and (3) it doesn’t contain irrelevant or extra information in addition to the justification itself.
      • Choose ’No’: otherwise.
    • Completeness:
      • Choose ’Yes’: A claim for completeness is correct if it follows ALL the three criteria: (1) it is a correct judgement, (2) it correctly justifies that the answer either covers the desired content to understand the result or points out the missing information that can make the result confusing or misleading, and (3) it doesn’t contain irrelevant or extra information in addition to the justification itself.
      • Choose ’No’: otherwise.
  2. 2.
    [Only proceed to this step if ’Claim_correct’ = ’Yes’] Evaluate the quality of the ’Citations’ generated by the evaluator LLM, and correct the citations if needed.
    For each claim, there’s a list of citations consisting of verbatim sentences extracted from ’Query’, ’Context’, and ’Answer’, with source specified in the parentheses, as evidence to support the claim. Please evaluate the correctness of the generated citations as below:
    • Choose ’Yes’ for ’Citations_correct’: The citations are correct if it follows ALL the three criteria: (1) the sources in the parentheses are correct, (2) no sentences are rewritten/paraphrased, and (3) they are all relevant evidence to support the claim, and there are no missing relevant sentences not cited from either ’Query’, ’Context’, or ’Answer’. You don’t have to do anything for ’Revised_citations’ in this case.
    • Choose ’No’ for ’Citations_correct’: If any of the three criteria is not met, choose ’No’ for ’Citations_correct’, and revise the generated citations to provide a correct list of citations (following the generated format) in ’Revised_citations’.
Appendix E Prompt for LLM-AggreFact Eval
Determine whether the provided claim is consistentwith the corresponding document. Consistency inthis context implies that all information presentedin the claim is substantiated by the document.If not, it should be considered inconsistent.Document: {doc}Claim: {claim}Please assess the claim’s consistency with thedocument by responding with either "yes" or "no".Answer (yes|no only):